estimate the doubling time of this function:
p= 2.1(1.0475)^ t where t is in years

Answer by stanbon(75887)

p= 2.1(1.0475)^ t where t is in years
If I read you correctly, 2.1 is the amount invested.
Double that would be 4.2.
EQUATION:
4.2 = 2.1(1.0475)^t
2=1.0475^t
Take the ln of both sides to get:
ln2 = t(ln1.0475)
t=ln2 / ln1.0475
t = 3.7635
----------
Cheers,
Stan H.